OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

If it's parallel, then it will have the same slope. After that, substitute in the x point and then change the b value until it gets the y value.

OpenStudy (anonymous):

so the slope would be -3 or 1/3?

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

If it's the same as -3, then it would be -3

OpenStudy (anonymous):

Ok and I the "b" value I got 3 but the choices give 7 too?

OpenStudy (anonymous):

What are your choices? And 3 would be the correct b value.

OpenStudy (anonymous):

y = 1/3x + 7 y= -3x + 3 y = 1/3x + 3 y = -3x + 7

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

@200205650

OpenStudy (anonymous):

Okay hold on. I'm just going to do this on my calculator.

OpenStudy (anonymous):

y=-3x+3 should be the answer. Do you have a graphing calculator?

OpenStudy (anonymous):

No but I can search one up online

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome!